# Supercritical fluids as a distinct state of matter characterized by sub-short-range structural order

### v1

@@ Abstract
A supercritical fluid (SCF) – the state of matter at temperatures and pressures above the critical point – exhibits properties intermediate between those of a liquid and a gas. However, whether it constitutes a fundamentally distinct phase or merely a continuous extension of the liquid and gas states remains an open question. Here we show that a SCF is defined by sub-short-range (SSR) structural order in the spatial arrangement of particles, setting it apart from the gas (disordered), liquid (short-range ordered), and solid (long-range ordered) states. The SSR structural order can be characterized by a length scale effectively quantified by the number of observable peaks in the radial distribution function. This length grows from a minimum microscopic value, on the order of the inter-particle distance at the gas–SCF boundary, to a diverging value at the SCF–liquid boundary. Based on the emergence of SSR order, we demonstrate that the transport and dynamical properties of the SCF state, including the diffusion coefficient, shear viscosity, and velocity autocorrelation function, also clearly distinguish it from both the liquid and gas states. Theoretical predictions are validated by molecular dynamics simulations of argon and further supported by existing experimental evidence. Our study confirms and assigns physical significance to the refined phase diagram of matter in the supercritical region, consisting of three distinct states (gas, supercritical fluid, and liquid) separated by two crossover boundaries that follow universal scaling laws.
@@
Introduction. Liquids and gases are arguably among the most common phases of matter Tabor (1991) and have been studied and probed experimentally since centuries. At low enough pressure and temperature, they are separated by a coexistence line defined by a thermodynamic first order phase transition Domb (2000). Nevertheless, the coexistence line has finite extension and ends into the so called critical point (e.g., $T_{c}\approx 647$ K and $P_{c}\approx 220$ bar for water), above which lies a large part of the phase diagram – the supercritical fluid (SCF) state of matter Kiran et al. (2012); Proctor (2020) – which was first discovered by Baron de la Tour in 1822 de La Tour (1822). In this SCF state, liquid and gas phases can be connected without encountering any thermodynamic singularity, which raises the long-standing question of what really separates a liquid from a gas Brazhkin and Trachenko (2012), if anything.
@@
A commonly considered scenario is that the SCF state is formed by liquid-like and gas-like states, which are separated in the phase diagram by a single separation line (e.g., Widom line, Frenkel line and Widom-Fisher line, etc.) Xu et al. (2005); Jones and Walker (1956); Luo et al. (2014); Banuti et al. (2017); Corradini et al. (2015); Gallo et al. (2014); Fisher and Widom (1969); Vega et al. (1995); Evans et al. (1993); Brazhkin et al. (2012a, 2013); Bolmatov et al. (2013, 2014); Fomin et al. (2018); Nishikawa and Tanaka (1995); Matsugami et al. (2014); Ploetz and Smith (2019); Woodcock (2013, 2016); Brazhkin et al. (2012b); Huang et al. (2023); Ouyang et al. (2024). An alternative scenario is that liquid and solid phases are separated by a crossover region, bounded by two boundary lines Banuti (2015); Maxim et al. (2021); Wang et al. (2021); Li and Jin (2024), making the SCF in this region a distinguishable state from a liquid or a gas. Recent studies have established the thermodynamic origin of these two supercritical boundary lines ($L^{\pm}$ lines)  Li and Jin (2024) and their universal scalings in liquid-gas phase transitions Li and Jin (2024), Ising models Li and Jin (2024), quantum phase transitions Wang et al. (2024); Lv et al. (2024), charged black holes Wang et al. (2025) and liquid-liquid phase transitions Wang et al. (2025).
@@
Figure 1: Phase diagram and radial distribution functions. (a) Phase diagram of argon. The critical point (CP) is located at $(T_{c},P_{c})\approx(151\,\text{K},48.55\,\text{bar})$ Mccain and Ziegler (1967). The liquid-gas coexistence line (solid black line) is obtained from the NIST database NIS . The gray dotted line is the critical isochore, which represents supercritical continuation of the coexistence line. The two black dashed lines are $L^{\pm}$ lines, which are computed using the NIST data and the thermodynamic criterion proposed in Li and Jin (2024), and represent the boundaries between liquid, gas and SCF phases. (b-d) Typical reduced RDF, $\tilde{h}(r)=r[g(r)-1]$, of gas, SCF and liquid states measured in simulations of argon (gas: $T=195$ K, $P=20~\text{bar}$, SCF: $T=195$ K, $90~\text{bar}$, liquid: $T=100$ K, $P=40~\text{bar}$).
If the SCF is an independent and separate state of matter as in the second scenario, complementary to the common liquid, gas and solid states, naturally one would ask what defines this state. Here we address this question focusing on structure, i.e., the spatial arrangement of particles. Our understanding is summarized in Fig. 1 by the phase diagram, and the behavior of a key structural indicator, the radial distribution function (RDF) $g(r)$ that describes the probability of finding a particle at a given distance $r$ from a reference particle. According to textbook knowledge, the structures of gases, liquids and solids (crystals) are characterized by disorder, short-range and long-range order, respectively. These structural features are captured by the behavior of $g(r)$ beyond the first peak, corresponding respectively to, exponential decay, exponentially damped oscillations, and undamped oscillations.
Here we propose that the structure of SCF is characterized by sub-short-range (SSR) order. The SSR order is reflected in $g(r)$ through damped oscillations that are truncated, i.e., the oscillatory behavior of $g(r)$ extends only up to a finite SSR length scale $\xi$. The length scale $\xi$ is ill-defined in the liquid phase, diverges at the liquid-SCF boundary ($L^{+}$ line), and reduces to the scale of nearest neighbor distance at the gas-SCF boundary ($L^{-}$ line).
This viewpoint on structure and, as we will see, also the behavior of transport and dynamical properties, imply that the SCF can be interpreted as an independent state of matter in a loose sense. This definition is not rigorous because there are no bona fide phase transitions between SCF and gases/liquids and all lines should be rather regarded as smooth crossovers. Nevertheless, these three phases can be still separated by two supercritical boundary lines:
(i) Liquid-SCF boundary (LSB). In our framework, the Fisher-Widom line Fisher and Widom (1969) ($\xi\to\infty$) corresponds to the liquid-SCF boundary. In liquids, the oscillations in $g(r)$, $\sim\cos(qr)$, although damped, remain up to $r\to\infty$. The LSB has been alternatively estimated based on dynamic criteria (Frenkel line Brazhkin et al. (2012a)), such as the disappearance of the short-time vibrational motion typical of liquids. The Fisher-Widom line and the Frenkel line are also close to the upper $L^{+}$ line defined from thermodynamic and scaling considerations Li and Jin (2024). In general, the liquid-SCF boundary corresponds to the boundary beyond which the state loses liquid characteristics, although its precise location depends on the detailed criteria. The small quantitative inconsistency is possibly inevitable (also for the gas-SCF boundary), due to the absence of thermodynamic or dynamic singularities. Nevertheless, these lines are close to each other, and can be unambiguously distinguished from the gas-SCF boundary (as we will show in this study). Since the liquid–SCF transition has been extensively studied in the literature, including from the perspective of structural criteria such as the Fisher-Widom line Fisher and Widom (1969); De Carvalho et al. (1994); Vega et al. (1995); Dijkstra and Evans (2000a); Evans et al. (1993); Tarazona et al. (2003); Stopper et al. (2019), we do not focus on it here. Instead, our study concentrates on the gas-SCF boundary, which remains largely unexplored.
(ii) Gas-SCF boundary. This boundary corresponds to the locus where the oscillatory behavior in $g(r)$, aside from the first peak which persists except in the ideal gas limit, disappears. In practice, it can be located by the disappearance of the second peak in $g(r)$, or equivalently the first minimum, as shown by the transition from Fig.1(b) to Fig.1(c), using molecular dynamics (MD) simulation data of argon (Ar). Below, we reveal that this structural change in $g(r)$ is reflected in transport properties, specifically, the viscosity $\eta$ and the self-diffusion coefficient $D$, as well as in the dynamical behavior, such as the velocity autocorrelation function (VACF), $Z(t)$. These observations can also be derived from a liquid-state theory based on the memory-function approach and are supported by simulation results. The numerically estimated gas–SCF boundary based on these criteria aligns closely with the lower $L^{-}$ line independently identified in Ref. Li and Jin (2024) using a thermodynamic criterion, suggesting a clear physical significance for the latter.
@@
As shown in Fig. 1a, we propose that the gas, SCF and liquid phases are separated by two boundary lines, (black dashed lines). (i) The liquid-SCF boundary corresponds to the Fisher-Widom line Vega et al. (1995) defined based on the behavior of $g(r)$ (a structural viewpoint), or the Frenkel line defined based on dynamical viewpoints. The location of the Frenkel line depends on the specific criterion employed: the Frenkel-1 line is defined by the equality $\tau_{0}=\tau$ between the vibration time $\tau_{0}$ and the liquid relaxation time $\tau$ Brazhkin et al. (2012a), the Frenkel-2 line corresponds to the disappearance of oscillations and minima of the velocity autocorrelation function Brazhkin et al. (2013), and the Frenkel-3 line is derived based on isochoric heat capacity $C_{V}$ (for monatomic systems such as argon, the criterion is $C_{V}=2k_{\rm B}$Brazhkin et al. (2012a). (ii) The gas-SCF boundary, which is the focus of the present study, has never been individually investigated in previous works. (iii) Several existing considerations have proposed two boundaries, including the $L^{\pm}$ lines Li and Jin (2024), percolation lines Woodcock (2013, 2016), and the two boundaries of the pseudo-boiling transitional region Banuti (2015); Wang et al. (2021). Different criteria give nearby, but quantitatively different boundary lines (see Li and Jin (2024)). Again, the reason is that these lines do not correspond to any thermodynamic or dynamic singularities, which means that they cannot be uniquely defined in a mathematically rigorous way even in the thermodynamic limit. An exception is the $L^{\pm}$ lines in quantum phase transitions, where the thermodynamic supercritical crossovers coincide with dynamic singularities under quantum quenches Wang et al. (2024).
@@
Among the above proposals, only the $L^{\pm}$ lines display universal scalings in various phase transitions Li and Jin (2024); Wang et al. (2024); Lv et al. (2024); Wang et al. (2025): the scalings of the external field and order parameter along $L^{\pm}$ lines satisfy $\delta P^{\pm}\sim(T-T_{\rm c})^{\beta+\gamma}$ and $\delta\rho^{\pm}\sim(T-T_{\rm c})^{\beta}$, where $\beta$ and $\gamma$ are standard critical exponents. Below we will use $L^{\pm}$ lines as a reference of the two boundaries (see Methods).
@@
For the detailed analyses below, we select two isothermal lines corresponding to $T=195$ K (results in the main text) and $T=220$ K (see Supplementary Material) respectively and consider a large range of pressures that scan from largely above the liquid-SCF boundary to well below the gas-SCF boundary. This will allow us to study in detail the structural, transport and dynamical properties across the boundaries, revealing their fundamental origin and nature. In particular, in the next section we will show that the emergence of SSR order in $g(r)$ coincides with the $L^{-}$ line.
Theory. From the Ornstein–Zernike equation Hansen and McDonald (2013), the asymptotic behavior of the reduced total correlation function $\tilde{h}(r)=r[g(r)-1]$ can be expressed in Fourier representation as $\tilde{h}(r)=\sum_{n}A_{n}e^{ik_{n}r}$. The poles $k_{n}$ satisfy the condition $1-\rho\,\hat{c}(k)=0$, where $\rho$ is the number density and $\hat{c}(k)$ is the Fourier transform of the pair direct correlation function $c(r)$. The coefficient $2\pi A_{n}$ represents the residue associated with the $n$-th pole.
In general, an infinite number of poles $k_{n}$ are expected. We emphasize that at least one of these two leading poles must be purely imaginary de Carvalho et al. (1994), while the others consist of complex-conjugate pairs with identical imaginary parts and real parts of opposite sign. However, several studies (e.g., Dijkstra and Evans (2000b)) have shown that retaining only the two leading poles, i.e., those with the smallest imaginary parts, provides an accurate description of $\tilde{h}(r)$ up to interatomic distances of approximately $2\sigma_{0}$, corresponding to second-nearest neighbors ($\sigma_{0}$ is defined by the location of the minimum in the inter-particle potential $V(r)$). Remarkably, this approximation also captures the correct position of the first peak in $\tilde{h}(r)$. This success was attributed to the fact that the next leading poles have a considerably larger imaginary part in a variety of systems de Carvalho et al. (1994); Leote de Carvalho et al. (1999); de Carvalho and Evans (1994); Carvalho and Evans (1997).
Supported by these observations, we then write $\tilde{h}(r)$ as the sum of two dominating terms, $\tilde{h}(r)=A_{0}e^{-r/r_{0}}+A_{1}e^{-r/r_{1}}\cos[q(r-\sigma_{0})]=A_{0}e^{-r/r_{0}}\{1+ae^{-r/\xi}\cos[q(r-\sigma_{0})]\}$, where $a=A_{1}/A_{0}$ and $1/r_{1}=1/r_{0}+1/\xi$ Dijkstra and Evans (2000a). As a direct numerical investigation will confirm, the wave vector $q\approx 2\pi/\sigma_{0}$ is nearly a constant in the interested regime. The first exponential term $e^{-r/r_{0}}$ and the second oscillating term $e^{-r/r_{1}}\cos[q(r-\sigma_{0})]$ correspond respectively to the leading pure imaginary and complex poles. In this language, the Fisher-Widom line can be elegantly defined by the condition $r_{1}=r_{0}$, or equivalently $\xi=\infty$  Fisher and Widom (1969). More precisely, in the liquid phase, $r_{1}>r_{0}$ and the asymptotic decay of $\tilde{h}(r)$ at the limit $r\to\infty$ is dominated by the oscillating term; in the gas phase, $r_{1}<r_{0}$ and the same asymptotic decay is dominated by the exponential term.
Using the two-pole decomposition described above and the newly introduced SSR length scale $\xi$, we propose structural criteria for both the liquid–SCF and gas–SCF boundaries (see Fig. 1(a)). These boundaries were previously identified through thermodynamic analysis Li and Jin (2024), but their precise physical interpretation has remained unclear. Within our proposal, the liquid-SCF boundary is determined by $\xi\to\infty$, which is identical to Fisher-Widom’s condition $r_{1}=r_{0}$ Fisher and Widom (1969); the gas-SCF boundary corresponds to $\xi=c_{0}\sigma_{0}$ where $c_{0}$ is an order one dimensionless constant. From a physical perspective, $\xi$ characterizes the degree of structural correlations in the system. In particular, $\xi=\infty$ corresponds to a strongly correlated many-body system, i.e., a liquid. In contrast, $\xi\sim\mathcal{O}(\sigma_{0})$ indicates a correlation length on the order of the nearest-neighbor distance, as expected in a gas described by a Lennard-Jones potential (Fig. 1(b)), where short-range order is totally lost.
A simple, though approximate, interpretation of $\xi$ is that $n=\xi/\sigma_{0}$ gives the number of peaks observable in $g(r)$. Given the simple expression $1+ae^{-r/\xi}\cos[q(r-\sigma_{0})]$ appearing in $\tilde{h}$, the $n$-th peak at $r=n\sigma_{0}$ is negligible when the second term $ae^{-r/\xi}\cos[q(r-\sigma_{0})]$ is much smaller than one, or $e^{-n\sigma_{0}/\xi}\ll 1$. In this sense, the gas-SCF boundaries correspond to $g(r)$ displaying only a single peak, followed by an exponential decay, as shown in Fig. 1(b). This argument gives a way to approximately estimate $\xi(P)$: as $P$ increases from the dilute limit, when the $n$-the peak appears at $P=P_{n}$, the length scale is evaluated as $\xi(P=P_{n})=n\sigma_{0}$ at $P_{n}$. This idea will guide our data analysis below.
The general picture we expect is the following. (i) In the gas state, the oscillation in $\tilde{h}(r)$ damps rapidly and disappears after the first peak $r\approx\sigma_{0}$ ($\xi/\sigma_{0}\approx n_{0}$; see Fig. 1b). This means that the spatial correlations between particles are negligible beyond a distance comparable to the nearest neighbor distance. The large distance decay of these correlations is determined by the length scale $r_{0}$. (ii) In the SCF state, the SSR structural order gradually grows from the gas-SCF to liquid-SCF boundaries, reflected by the increase of the number of peaks in $\tilde{h}(r)$ from $\xi/\sigma_{0}\approx n_{0}$ to $\xi/\sigma_{0}\to\infty$. In the SCF state, the oscillations are damped and survive up to a finite distance $\xi$ (see Fig. 1c). This corresponds to the gradual formation of higher-order coordination shells around a central tagged particle in the SCF state. (iii) In the liquid state, the oscillations are still damped but extend to an infinitely large distance (see Fig. 1d), as originally advocated by Fisher and Widom Fisher and Widom (1969). The asymptotic decay of these oscillations in the liquid states is characterized by the length scale $r_{1}$.
Although the decomposition in terms of the two length scales $r_{0}$ and $r_{1}$ is mathematically more straightforward, we find that the SSR order length scale, defined as $\xi\equiv r_{0}r_{1}/(r_{0}-r_{1})$, offers a more physically transparent interpretation within the SCF phase. Importantly, all of our conclusions can be equivalently expressed in terms of $r_{0}$ and $r_{1}$, which remain well-defined across all three phases, up to the emergence of long-range order in ideal crystalline states, where they diverge. However, our analysis suggests that reliably extracting $r_{0}$ and $r_{1}$ by directly fitting $g(r)$ requires data of very high precision, making this procedure quite challenging in practice.
@@
Figure 2: SSR structural order and characteristic lengths. (a) Radial distribution function $g(r)$ at different pressures, zoomed in around the second peak, for T=195KT=195\ $\text${K}. (inset) Height of the second peak $h_{2}$ as a function of $P$; below $P_{2}\approx 55$ bar (red circle), the second peak can not be identified anymore. (b) The SSR length $\xi$ and asymptotic decay length $r^{*}$ as functions of $P$ along the isothermal line T=195KT=195\ $\text${K}.

### v2

# Physical Signatures of Supercritical Fluid Boundaries

@@ Abstract
In the supercritical fluid (SCF) region, at temperatures and pressures above the critical point, the thermodynamic singularity separating liquids and gases no longer exists. Recent arguments based on thermodynamics and critical scalings have revived the proposal that the SCF constitutes an intermediate state of matter, separated from the liquid and gas by two supercritical boundaries, the $L^{\pm}$ lines. However, until now, the nature of the supercritical state and the physical signatures of these boundaries have remained elusive. Here, we demonstrate that the SCF is characterized by distinct structural, transport, and dynamical behavior. Specifically, the spatial arrangement of particles—captured by the radial distribution function—as well as the diffusion coefficient, shear viscosity, and velocity autocorrelation function in the SCF regime are qualitatively different from those in both the liquid and gas states and exhibit clear physical signatures upon crossing the $L^{\pm}$ lines. Our theoretical predictions are validated by molecular dynamics simulations of argon and are further supported by existing experimental evidence. These results provide a clear physical foundation for a refined phase diagram of matter in the supercritical region, comprising three distinct states—gas, supercritical fluid, and liquid—separated by two crossover boundaries obeying universal scaling laws.
@@
Introduction. Liquids and gases are arguably among the most common phases of matter Tabor (1991) and have been studied and probed experimentally since centuries. At low enough pressure and temperature, they are separated by a coexistence line defined by a thermodynamic first order phase transition Domb (2000). Nevertheless, the coexistence line has finite extension and ends into the so called critical point (e.g., $T_{c}\approx 647$ K and $P_{c}\approx 220$ bar for water), above which lies a large part of the phase diagram – the supercritical fluid (SCF) region Kiran et al. (2012); Proctor (2020) – which was first discovered by Baron de la Tour in 1822 de La Tour (1822). In this SCF region, liquid and gas phases can be connected without encountering any thermodynamic singularity, which raises the long-standing question of what really separates a liquid from a gas Brazhkin and Trachenko (2012), if anything, and whether a supercritical fluid constitutes a distinct phase of matter.
@@
A commonly considered scenario is that the SCF region contains liquid-like and gas-like states separated in the phase diagram by a single separation line. Importantly, this line does not correspond to any thermodynamic singularity and therefore represents a continuous crossover, making both its definition and precise location inherently subtle. Indeed, multiple distinct and non-coincident definitions of this crossover line have been proposed in the literature, including the Widom line, Frenkel line, and Widom–Fisher line, among others Xu et al. (2005); Jones and Walker (1956); Luo et al. (2014); Banuti et al. (2017); Corradini et al. (2015); Gallo et al. (2014); Fisher and Widom (1969); Vega et al. (1995); Evans et al. (1993); Brazhkin et al. (2012a, 2013); Bolmatov et al. (2013, 2014); Fomin et al. (2018); Nishikawa and Tanaka (1995); Matsugami et al. (2014); Ploetz and Smith (2019); Woodcock (2013, 2016); Brazhkin et al. (2012b); Huang et al. (2023); Ouyang et al. (2024).
An alternative scenario, summarized in Fig. 1, posits that the liquid and gas phases are separated by a finite crossover region, bounded by two boundary lines Banuti (2015); Maxim et al. (2021); Wang et al. (2021); Li and Jin (2024), thereby rendering the SCF in this region distinguishable from both a liquid and a gas. A two-boundary picture has been explored, for instance, in the Widom delta scenario Ha et al. (2018); Yoon et al. (2019); Ha et al. (2020), where the crossover lines are associated with the percolation loci of gas-like and liquid-like clusters. In this framework, however, the intermediate SCF region is not treated as a distinct state of matter but rather as a gas–liquid coexistence regime. Moreover, the definition of such clusters is often ambiguous and typically relies on machine-learning-based classifications Ha et al. (2018), lacking a definitive interpretation of the associated physical signatures (see however Asharchuk et al. (2025)).
On the other hand, more recent studies have established the thermodynamic origin of these two supercritical boundary lines ($L^{\pm}$ lines)  Li and Jin (2024) and their universal scalings in liquid-gas phase transitions Li and Jin (2024), Ising models Li and Jin (2024), quantum phase transitions Wang et al. (2024); Lv et al. (2024), charged black holes Wang et al. (2025) and liquid-liquid phase transitions Wang et al. (2025).
If, however, the SCF constitutes an independent state of matter, as implied by the second scenario Li and Jin (2024), then a natural question arises: what defines this state, and what physical signatures mark the crossing of the two supercritical boundaries? Here, we address this question by showing that the SCF exhibits unique structural, transport, and dynamical properties, fundamentally different from those of both liquids and gases. These differences produce clear, albeit non-critical, physical signatures when crossing each of the two supercritical boundaries, consistently observed in both simulations and experiments.
@@
Figure 1: Phase diagram of argon. The critical point (CP) is located at $(T_{c},P_{c})\approx(151\,\text{K},48.55\,\text{bar})$ Mccain and Ziegler (1967). The liquid-gas coexistence line (solid black line) is obtained from the NIST database NIS . The gray dotted line is the critical isochore, which represents supercritical continuation of the coexistence line. The two black dashed lines are $L^{\pm}$ lines, which are computed using the NIST data and the thermodynamic criterion proposed in Li and Jin (2024), and represent the boundaries between liquid, gas and SCF phases.
@@
As shown in Fig. 1a, we propose that the gas, SCF and liquid phases are separated by two boundary lines, (black dashed lines). (i) The liquid-SCF boundary roughly corresponds to the Fisher-Widom line Vega et al. (1995), defined based on the behavior of the radial distribution function $g(r)$ (a structural viewpoint), or the Frenkel line defined based on dynamical viewpoints. The location of the Frenkel line depends on the specific criterion employed: the Frenkel-1 line is defined by the equality $\tau_{0}=\tau$ between the vibration time $\tau_{0}$ and the liquid relaxation time $\tau$ Brazhkin et al. (2012a), the Frenkel-2 line corresponds to the disappearance of oscillations and minima of the velocity autocorrelation function Brazhkin et al. (2013), and the Frenkel-3 line is derived based on isochoric heat capacity $C_{V}$ (for monatomic systems such as argon, the criterion is $C_{V}=2k_{\rm B}$Brazhkin et al. (2012a). (ii) The gas-SCF boundary, which is the focus of the present study, has never been individually investigated in previous works. (iii) Several existing considerations have proposed two boundaries, including the $L^{\pm}$ lines Li and Jin (2024), percolation lines Woodcock (2013, 2016), the two boundaries of the pseudo-boiling transitional region Banuti (2015); Wang et al. (2021), and the Widom delta Ha et al. (2018); Yoon et al. (2019); Ha et al. (2020). Different criteria give nearby, but quantitatively different boundary lines (see Li and Jin (2024)). Again, the reason is that these lines do not correspond to any thermodynamic or dynamic singularities, which means that they cannot be uniquely defined in a mathematically rigorous way even in the thermodynamic limit. An exception is the $L^{\pm}$ lines in quantum phase transitions, where the thermodynamic supercritical crossovers coincide with dynamic singularities under quantum quenches Wang et al. (2024).
@@
Importantly, among the above proposals, only the $L^{\pm}$ lines display universal scalings in various phase transitions Li and Jin (2024); Wang et al. (2024); Lv et al. (2024); Wang et al. (2025): the scalings of the external field and order parameter along $L^{\pm}$ lines satisfy $\delta P^{\pm}\sim(T-T_{\rm c})^{\beta+\gamma}$ and $\delta\rho^{\pm}\sim(T-T_{\rm c})^{\beta}$, where $\beta$ and $\gamma$ are standard critical exponents. Because of these reasons, below, we adopt the $L^{\pm}$ lines, introduced in Ref. Li and Jin (2024) and shown for the case of argon as black dashed curves in Fig. 1, as the reference for the two supercritical boundaries (see Methods).
@@
For the detailed analyses below, we select two isothermal lines corresponding to $T=195$ K (results in the main text) and $T=220$ K (see Supplementary Material Sec. S1) respectively and consider a large range of pressures that scan from largely above the liquid-SCF boundary to well below the gas-SCF boundary. This will allow us to study in detail the structural, transport and dynamical properties across the boundaries, revealing their fundamental origin and nature.
@@
Figure 2: Radial distribution functions (RDFs). (a-c) Typical reduced RDF, $\tilde{h}(r)=r[g(r)-1]$, of gas, SCF and liquid states measured in simulations of argon (gas: $T=195$ K, $P=20~\text{bar}$; SCF: $T=195$ K, $90~\text{bar}$; liquid: $T=100$ K, $P=40~\text{bar}$). The distance $r$ is expressed in units of $\sigma_{0}$, the position of the first peak. (d) Radial distribution function $g(r)$ at different pressures, zoomed in around the second peak, for T=195KT=195\ $\text${K}. (inset) Height of the second peak $h_{2}$ as a function of $P$; below $P_{2}\approx 55$ bar (red circle), the second peak can not be identified anymore.
Structural evidence: radial distribution functions. In Fig. 2(a-c), we present the reduced radial distribution function, $\tilde{h}(r)=r[g(r)-1]$, for argon in three distinct states: gas, SCF, and liquid. The gas state exhibits a rapid decay of the tail of $\tilde{h}(r)$ without oscillations, indicating the weak intermolecular correlations. In contrast, the liquid displays exponentially damped (yet persistent) oscillations with multiple distinct peaks, signaling robust short-range order. The SCF exhibits an intermediate structure: a few peaks emerge, yet higher-order coordination shells are absent, indicating loss of orders at large distances. This represents a unique structural signature of the SCF, setting it apart from the gas and liquid states.
Based on above evidence, we propose that the structure of SCF is characterized by sub-short-range (SSR) order. The SSR order is reflected in the radial distribution function $g(r)$ through damped oscillations that are truncated, i.e., the oscillatory behavior of $g(r)$ extends only up to a finite SSR length scale. This viewpoint on structure and, as we will see, also the behavior of transport and dynamical properties, imply that the SCF can be interpreted as an independent state of matter in a loose sense. This definition is not rigorous because there are no bona fide phase transitions between SCF and gases/liquids and all lines should be rather regarded as smooth crossovers. Nevertheless, these three phases can be still separated by two supercritical boundary lines:
(i) Liquid-SCF boundary. From a structural perspective, the Fisher-Widom line Fisher and Widom (1969) corresponds to the liquid-SCF boundary. In liquids, the oscillations in $g(r)$, $\sim\cos(qr)$, although damped, remain up to $r\to\infty$. The liquid-SCF boundary has been alternatively estimated based on dynamic criteria (Frenkel line Brazhkin et al. (2012a)), such as the disappearance of the short-time vibrational motion typical of liquids. The Fisher-Widom line and the Frenkel line are also close to the upper $L^{+}$ line defined from thermodynamic and scaling considerations Li and Jin (2024). In general, the liquid-SCF boundary corresponds to the boundary beyond which the state loses liquid characteristics, although its precise location depends on the detailed criteria. The small quantitative inconsistency is possibly inevitable (also for the gas-SCF boundary), due to the absence of thermodynamic or dynamic singularities. Nevertheless, these lines are close to each other, and can be unambiguously distinguished from the gas-SCF boundary (as we will show in this study). Since the liquid–SCF transition has been extensively studied in the literature, including from the perspective of structural criteria such as the Fisher-Widom line Fisher and Widom (1969); De Carvalho et al. (1994); Vega et al. (1995); Dijkstra and Evans (2000); Evans et al. (1993); Tarazona et al. (2003); Stopper et al. (2019), we do not focus on it here. Instead, our study concentrates on the gas-SCF boundary, which remains largely unexplored.
(ii) Gas-SCF boundary. This boundary corresponds to the locus where the oscillatory behavior in $g(r)$, aside from the first peak which persists except in the ideal gas limit, disappears. In practice, it can be located by the disappearance of the second peak in $g(r)$, or equivalently the first minimum, as shown by the transition from Fig.1(b) to Fig.1(c), using MD simulation data of argon.